OpenStudy (anonymous):

Find the domain of the function: g(x) = 4/8-9x

OpenStudy (anonymous):

Hey by budster!!!

OpenStudy (dumbcow):

Domain is either All real numbers or x does not equal 8/9 can you clarify the function, whats in denominator?

OpenStudy (saifoo.khan):

OOOO Heyyy! i didnt notice thts u, coz u chgnd your photo! How are u? sorry i left that day coz there was electric failure!

OpenStudy (anonymous):

Is the demonintaor the one one the top of on the bottom? I can't remember.

OpenStudy (saifoo.khan):

denom is the down one.

OpenStudy (anonymous):

No problem...

OpenStudy (anonymous):

Okay, that is a good way to remember...d for down.

OpenStudy (dumbcow):

is it...4/(8-9x) or (4/8) - 9x

OpenStudy (saifoo.khan):

Oh hmm.

OpenStudy (anonymous):

no cow, those are not any of my options.

OpenStudy (dumbcow):

??? what is g(x)? 4/(8-9x) or (4/8) - 9x

OpenStudy (anonymous):

Okay, this is a multiple question. Here are my choices: x is a real number and x is not equal to 4 x is a real number is and not equal to 8/9 x is a real number and is >= to 8/9 x is a real number is is <= 0

OpenStudy (zarkon):

dumbcow already gave you the answer

OpenStudy (anonymous):

Yea Safoo, I finally decided to change my avatar. I got tired of being an "owl." LOL

OpenStudy (anonymous):

I'm sorry I didn't see it.

OpenStudy (saifoo.khan):

Haha, cool.

OpenStudy (dumbcow):

2nd one set denominator equal to zero 8-9x = 0 solve for x -9x = -8 x = -8/-9 = 8/9 so at x=8/9, g is undefined so it cant be part of the domain

OpenStudy (anonymous):

Thank you dumbcow.

OpenStudy (anonymous):

damn it. So can you please say it in plain english? Is it less than, more than or equal to 8/9? Or something else?

OpenStudy (anonymous):

No disrespect meant...that didn't really come out how I meant it.

OpenStudy (saifoo.khan):

LOL.

OpenStudy (anonymous):

I think I made them mad. that really didn't come out how I meant for it to.

OpenStudy (anonymous):

So.....does anyone have the answer?

OpenStudy (saifoo.khan):

haha, wait let me figure it out.

OpenStudy (anonymous):

thanks hun....

OpenStudy (saifoo.khan):

the answer is second options! :]